OpenStudy (anonymous):

A blue bike is $14 less than a red bike. The sum of their prices is $300. How much is the red bike? The red bike is $____. i thought it would be 164 but apparently that's wrong.

OpenStudy (anonymous):

Let x = cost of blue bike x+14 = cost of a red bik x + x + 14 = 300 2x + 14 = 300 2x = 286 x = $143 therefore blue bike is $143 and red bike is $143+14 = $157 If you add those together, it comes up to $300

OpenStudy (anonymous):

smarty pants

OpenStudy (anonymous):

thanks lol but nah i just had this question like yesterday so i remembered the answer

OpenStudy (anonymous):

@skinny23 see i know a lil bit of somethin ;)